Calculate the ionic radii of Na+ and F- ions in NaF crystals ( inerionic distance of Na+ and F- is 2.31A°.)

Grade:12th pass

1 Answers

Arun
5 years ago
 
Dear Kajal
 
Radius of sodium ion =  2.31 * (4.5) /(6.5+ 4.5) = 0.95 A
similarly
 
radius of fluoride ion = 2.31 * (6.5) /(4.5+6.5) = 1.36 A
 
 
hope it helps
 
Regards
Arun (askIITians forum expert)